OpenStudy (anonymous):

P(t)=P(o)e raised to the r times t. P(o)=5,000 r=0.06 t=8. (a) quarterly. (b) continuously.

OpenStudy (anonymous):

here is the equation, \[P(t)=P(o)e ^{rt}\]

OpenStudy (anonymous):

wat is the different between finding it quarterly and finding it continuously?

OpenStudy (anonymous):

sarahk2, do you think you could help me out with this problem?

OpenStudy (anonymous):

kc, any tips?

OpenStudy (anonymous):

you guys there?

OpenStudy (anonymous):

hello?

OpenStudy (kc_kennylau):

I don't understand the question so many created symbols

OpenStudy (anonymous):

ahh

OpenStudy (kc_kennylau):

wait i actually get it

OpenStudy (kc_kennylau):

e is involved. Definitely continuously.

OpenStudy (anonymous):

so if I just plug in the values, I get 8,080.372011

OpenStudy (anonymous):

would that be for continuous or quarterly?

OpenStudy (anonymous):

you there?

OpenStudy (kc_kennylau):

If you see e in the formula then it's definitely continuously

OpenStudy (anonymous):

yeah, you are back :)

OpenStudy (anonymous):

so that is it continuously

OpenStudy (anonymous):

but how bout quarterly?

OpenStudy (anonymous):

do I just divide by 4 or use a different formula?

OpenStudy (kc_kennylau):

P(t)=P(o)(1+r/4)^(4t) would be quaterly

OpenStudy (anonymous):

Hero!!!!!!!

OpenStudy (anonymous):

wat up?

OpenStudy (anonymous):

so for quarterly

OpenStudy (anonymous):

let me get this straight

hero (hero):

Yeah, but...

OpenStudy (kc_kennylau):

sorry let me finish my breakfast first

OpenStudy (anonymous):

the formula would be for quarterly the following:? \[P(t)=p(o)(1+r/4)^{4t}\]

hero (hero):

There you go

OpenStudy (anonymous):

ok, so wat happened to the e?

hero (hero):

As @kc_kennylau was saying...if you see \(e\), in a compound interest formula, it is for continuous compounding

OpenStudy (anonymous):

by the way, I plugged in the values for quarterly and got appromixmately 8,051.6216

OpenStudy (anonymous):

but

OpenStudy (anonymous):

how do I derive those formulas for quarterly and continuous?

OpenStudy (anonymous):

will the professor have them on the test or is it just like remembering the pythagorean theorem?

OpenStudy (anonymous):

any comments, hero or kc?

hero (hero):

You derive them from Exponential equation: \[y = ab^x\] (x,y) = point on the curve a = the initial value b = growth or decay rate

OpenStudy (anonymous):

ty, hero and kc :)

OpenStudy (anonymous):

I really appreicate it

hero (hero):

(x,y) can refer to things such as (time, population) (years, account balance)